A traffic cone is 19 inches tall & has a radius of 5 inches. Find the LATERAL area.
Answer by Engg.N Magampon(6)

given:
h=19 inches
r=5 inches
LA=?
LA=pi*r*s
not knowing the value of S we will use pythagorean theorem
s^2=h^2+r^2
s^2=19^2 + 5^2
s^2= 361 + 25
s^2=386
s = square root of 386
s=19.86
LA=(3.1416)(5)(19.86)


LA=311.96